Girard wishes to mix red grapes worth $1.90 a pound with green grapes worth $1.10 a pound to get a 24 pound mixture worth $1.50 a pound. How many pounds of each does he need?
please help me

Answer by mananth(12270)

Girard wishes to mix red grapes worth $1.90 a pound with green grapes worth $1.10 a pound to get a 24 pound mixture worth $1.50 a pound. How many pounds of each does he need?
...
red grapes worth $1.90 a pound ................x pounds
green grapes worth $1.10 a pound ..............24-x
mixture 24 pound mixture worth $1.50 a pound.
..
1.90x+1.10(24-x)=24*1.50
1.90x+26.4-1.10x=36
0.8x+26.4=36
-26.4 -26.4
0.8x=9.6
/0.8
x=12 pounds of red grapes
12 pounds of green grapes
...
m.ananth@hotmail.ca